OpenStudy (anonymous):

Write an equation in point-slope form of the line that passes through the points (-5,-4) and (7,-5).

OpenStudy (anonymous):

@jim_thompson5910 @mathmath333

jimthompson5910 (jim_thompson5910):

same steps as before since they want it in point-slope form

OpenStudy (anonymous):

y-ysub1=m+(x-xsub1) right

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

you can just write y1 instead of "ysub1"

jimthompson5910 (jim_thompson5910):

same with x1

jimthompson5910 (jim_thompson5910):

oh it's not m+(x-x1)

OpenStudy (anonymous):

oh ok

jimthompson5910 (jim_thompson5910):

it's m times (x-x1)

OpenStudy (anonymous):

yeah i know so anyway \[m=\frac{ -5-(-4) }{ 7-(-5) }\]

OpenStudy (anonymous):

m=-1/12

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

y+4=-1/2(x+5) right

jimthompson5910 (jim_thompson5910):

I think you meant to say -1/12 instead of -1/2

OpenStudy (anonymous):

yeah i did

jimthompson5910 (jim_thompson5910):

its correct

OpenStudy (anonymous):

ok